OpenStudy (anonymous):

what is 4^3 (sqrt 135)????

OpenStudy (anonymous):

4^3 = 64; 135 = 5*27, which = 5*3*9, so the square root of 135 = 3*sqrt of 15. 3 *64 = 192, so the final answer is \[192\sqrt{15}\]